QUESTION 1

  1. Anglo Foods Inc., has the following forecasted sales for 2018:

    January

    February

    March

    April

    May

    June

    July

    $200,000

    $350,000

    $420,000

    $350,000

    $270,000

    $200,000

    $300,000

    Actual sales in November and December 2017 were $375,000 and 266,667, respectively. 60% of sales are on credit (i.e. 40% of sales will be received in cash in the same month). The firm collects 60% of these credit sales (i.e., 60%*60% = 36% of sales) during the first month after the sale and the remainder (i.e. 40% of credit sales and thus 40%*60% = 24% of sales) during the following second month.

    For example, for the January sales of $200,000, 40% (i.e. $80,000) will be received in cash in January, 36% (i.e. $72,000) will be received in cash in February, and the rest ($48,000) will be received in cash in March.

           Here are the information about the cash expenditures.

    1. Purchases constitute 60% of the next month’s sales. The company pays 50% during the first month after the purchase was made and the remainder in the next month.

    For example, January purchases will be 60% of February sales, i.e., $350,000*60%=$210,000. The company pays 50% of it ($105,000) in cash in February, and the rest in March.

    1. Wages, taxes and other expenses are expected to be 30% of forecasted sales.
    2. A major capital expenditure of $70,000 is expected in March.
    3. Interest payments are expected to be $30,000 every month.

    The company needs a minimum cash balance of $25,000. The beginning cash balance is $30,000.

    Using the Bithlo Barbecues’ simple cash budget example from Chapter 4, prepare their cash budget for January to June 2018.

    Answer Q (1-5) based on the cash budget you have prepared.

QUESTION 2

  1. How much is the total cash expenditure (total disbursement) in January?

    $230,00

    $66,000

    $266,000

    $210,000

7 points   

QUESTION 3

  1. How much is the unadjusted cash balance for March?

    $98,000

    ($73,000)

    $42,000

    $45,000

7 points   

QUESTION 4

  1. How much is the cash collection for January that is from sales one month earlier (December 2017)?

    $90,000

    $64,000

    $72,000

    $96,000

7 points   

QUESTION 5

  1. How much is the cumulative borrowing in March?

    $90,000

    $98,000

    $88,000

    $70,000
Answer #1
Cash Budget
January February March April May June
Beginning Cash Balance $         30,000 $         66,000 $         42,000 $        -73,000 $        -63,800 $        -26,000
Cash Collections $       266,000 $       276,000 $       342,000 $       375,200 $       334,800 $       261,200
Total Cash Available $       296,000 $       342,000 $       384,000 $       302,200 $       271,000 $       235,200
Cash Payments
Purchases $       140,000 $       165,000 $       231,000 $       231,000 $       186,000 $       141,000
Wages, taxes and other expenses $         60,000 $       105,000 $       126,000 $       105,000 $         81,000 $         60,000
Interest Payments $         30,000 $         30,000 $         30,000 $         30,000 $         30,000 $         30,000
Capital Expenditure $         70,000
Total Cash Payments $       230,000 $       300,000 $       457,000 $       366,000 $       297,000 $       231,000
Ending Cash balance $         66,000 $         42,000 $        -73,000 $        -63,800 $        -26,000 $            4,200

Working

Cash Budget
January February March April May June
Beginning Cash Balance
Cash Collections =200000*40%+266667*36%+375000*24% =350000*40%+200000*36%+266667*24% =420000*40%+350000*36%+200000*24% =350000*40%+420000*36%+350000*24% =270000*40%+350000*36%+420000*24% =200000*40%+270000*36%+350000*24%
Total Cash Available
Cash Payments
Purchases =266667*60%*50%+200000*60%*50% =200000*60%*50%+350000*60%*50% =350000*60%*50%+420000*60%*50% =420000*60%*50%+350000*60%*50% =270000*60%*50%+350000*60%*50% =200000*60%*50%+270000*60%*50%
Wages, taxes and other expenses =200000*30% =350000*30% =420000*30% =350000*30% =270000*30% =200000*30%

Q2.
Answer is a. $230,000

Q3.
Answer is b. ($73,000)

Q4.
Answer is d. $96000 i.e. 266667*36%

Q5.
Answer is b .$98000 i.e. $73000+25000